# Jinhong Li1    Qiuping Wang1    Shujie Han2    Patrick P. C. Lee1

### v2

@@ 3.2 Group-Based Data Layout
Trade-off analysis.
The choice of $G$ determines the trade-off between the degree of intra-zone parallelism via Zone Append and the stripe management overhead. A larger $G$ allows more stripes to be issued via Zone Append in parallel, but it also increases the stripe management overhead. In the extremes, when $G=1$, all stripes are issued via Zone Write; when $G=S$, there is only one stripe group in each segment (i.e., the group-based data layout is disabled for Zone Append).
@@
In contrast, the case of $G=S$ translates to 19 bits per stripe ID and hence 8.95 GiB of memory for the compact stripe table. Also, its high query overhead significantly increases the degraded read latency (§5.2).
Our prototype rounds up the length of a stripe ID to the nearest byte and represents a stripe ID in $\lceil\lceil\log_{2}G\rceil/8\rceil$ bytes. Representing stripe IDs in units of bytes, instead of bits, simplifies data alignment without compromising our conclusion.
@@ 3.3 Hybrid Data Management
Segment management.
Figure 5 shows how ZapRAID manages multiple open small-chunk and large-chunk segments. Let $N_{s}$ and $N_{l}$ be the numbers of open small-chunk segments and open large-chunk segments, respectively (i.e., the total number of open zones in each ZNS SSD is $N_{s}+N_{l}$). Also, let $C_{s}$ and $C_{l}$ be the sizes of a small chunk and a large chunk, respectively. All of $N_{s}$, $N_{l}$, $C_{s}$, and $C_{l}$ are configurable parameters. For example, we can set $C_{s}=8$ KiB for the small chunk size and $C_{l}=16$ KiB for the large chunk size. We can also set $N_{s}>1$ and $N_{l}>1$ to accommodate the real-world workloads that are mixed with small and large writes (e.g., in cloud block storage [55]).
@@
Choosing segments from in-flight stripes.
ZapRAID maintains in-flight stripes separately for small and large chunks. It assigns new written blocks to an in-flight stripe for either small or large chunks based on the write request size (see above). When an in-flight stripe is formed, ZapRAID forwards the in-flight stripe to an open segment. Specifically, if the in-flight stripe holds large chunks, ZapRAID selects an open large-chunk segment (out of the $N_{l}$) in a round-robin manner and writes the in-flight stripe to the selected segment via Zone Write. On the other hand, if the in-flight stripe holds small chunks, ZapRAID further considers two cases of $N_{s}$. If $N_{s}=1$, it writes the in-flight stripe to the only open small-chunk segment via Zone Append. If $N_{s}>1$, it selects an idle small-chunk segment (out of the $N_{s}-1$) in a round-robin manner and writes the in-flight stripe to the selected segment via Zone Write; however, if there is no idle small-chunk segment, ZapRAID selects the only small-chunk segment reserved for Zone Append and writes the in-flight stripe to the segment via Zone Append.
@@ 3.4 Crash Consistency
Recovery of the segment table.
ZapRAID first identifies the open and sealed segments in the system, so as to recover the segment table. Specifically, for each ZNS SSD, ZapRAID checks the states of all zones. If a zone is open or full, ZapRAID retrieves its first block, which belongs to the header region of a segment (§3.1). ZapRAID then reconstructs the segment table with the information of all possible segments based on the first blocks of all open and full zones.
@@
Recovery of stripes.
ZapRAID needs to recover the consistent states of all stripes by ensuring that each stripe has persisted $k$ data chunks and $m$ parity chunks. Specifically, ZapRAID examines whether each in-flight stripe has all its $k+m$ chunks persisted. Under the group-based data layout (note that $G=1$ means the use of Zone Write), ZapRAID ensures that only the stripes in the latest stripe group of an open segment can be the in-flight stripes. It retrieves all the chunks in the latest stripe group and examines the stripe IDs in their block metadata. ZapRAID discards all partially persisted stripes with fewer than $k+m$ persisted chunks. Recall that ZapRAID acknowledges the writes of an in-flight stripe only after persisting all $k+m$ chunks of the stripe (§3.1), so discarding the partially persisted stripes does not incur data loss. Since ZNS SSDs cannot perform in-place writes to replace the partially persisted chunks, ZapRAID rewrites all fully persisted stripes in the segment to a new segment and reclaims the old segment.
Recovery of the L2P table and the compact stripe table.
After ensuring the consistency of all stripes, ZapRAID recovers the L2P table and the compact stripe table (the latter is for segments that use Zone Append). Recall that for each sealed segment, ZapRAID stores the block metadata (i.e., the LBA, write timestamp, and stripe ID) of all blocks in the segment in the footer region, while for each open segment, ZapRAID also stores the block metadata in the out-of-band area associated with each block. To recover the L2P table and the compact stripe table, ZapRAID first retrieves the footer regions of all sealed segments and the data regions of all open segments. It examines all the block metadata and reconstructs the LBA-to-PBA mappings in the L2P table and the stripe IDs in the compact stripe table. For the L2P table, if multiple blocks have the same LBAs, ZapRAID keeps the one with the latest write timestamp.
@@ 3.5 Complete Workflows
Writes.
To issue a write request (identified by an LBA), ZapRAID checks whether the write request is a small or large one based on the write size, and writes accordingly the data to an in-flight stripe in memory for small chunks or large chunks (§3.3). If a new data chunk is available, ZapRAID issues a Zone Append or Zone Write command to write the chunk and block metadata based on the chunk’s position in the stripe and the RAID scheme. When an in-flight stripe contains $k$ data chunks, ZapRAID encodes them to generate $m$ parity chunks and their block metadata. It then issues a Zone Append or Zone Write command for each of the parity chunks to its respective zone. Only after all the chunks of a stripe are persisted, ZapRAID updates the L2P table with the corresponding LBAs and PBAs as well as acknowledges the completion of the write requests; the in-flight stripe is also released from memory. If there are insufficient data chunks to form a full stripe after a small timeout since the stripe is created (currently set as 100 $\mu$s in our prototype, which is close to the median write latency of 16-KiB chunks (§5.2)), ZapRAID fills the remaining stripe with zero blocks and invalid LBAs.
@@
Reads.
To read a block (identified by an LBA), ZapRAID queries the L2P table for its PBA (i.e., the segment ID, the drive ID, and the offset in the respective zone). Given the segment ID and drive ID, ZapRAID locates the zone from the segment table and retrieves the block from the specified offset of the zone.
Degraded reads.
Suppose that ZapRAID performs a degraded read to a lost block. Based on the PBA of the requested block, ZapRAID checks whether the segment applies Zone Append or Zone Write. For Zone Append, ZapRAID queries the compact stripe table to find out the stripe ID of the requested block. Since all available chunks of the same stripe must reside in the same stripe group under the group-based data layout, ZapRAID searches for the offsets of the $k$ available chunks in the stripe group from the compact stripe table; note that the search is efficient due to the limited group size (§5.2). It then reads the available chunks from the other zones, decodes the lost chunk, and retrieves the requested block. For Zone Write, ZapRAID reads the available chunks from other zones at the same offset to recover the chunk and hence the requested block.
Full-drive recovery.
When a drive fails, ZapRAID recovers the lost data into a new drive. It first identifies the segments that contain the lost zones in the failed drive by examining the segment-to-zones mappings in the header regions of all stored segments. For each lost zone, ZapRAID retrieves all available zones in the same segment from other available drives into memory, and reconstructs the stripe groups that cover the lost zone. It examines the block metadata of the available chunks and identifies the chunks from the same stripe. It then decodes the lost chunks for each stripe independently. After repairing all the stripes in a segment, ZapRAID writes the recovered zone to the new drive.
@@ 3.6 Discussion
Choice of the write request size threshold in hybrid data management.
Recall that ZapRAID compares the size of a write request against a threshold to determine if the write request is assigned to a small-chunk or large-chunk open segment (§3.3). Currently, the threshold is set to $C_{l}$. The choice of the threshold poses different performance trade-offs. A larger threshold implies that more writes will go to small-chunk segments. Full small-chunk stripes are formed faster, and hence small writes are also acknowledged faster (i.e., lower response time). However, the write throughput is lower since writing small-chunk segments is slower than writing large-chunk segments (based on the results in Figure 2 for different write request sizes). For simplicity, we assume that the threshold is equal to $C_{l}$ (i.e., if a write request size is smaller than $C_{l}$, it is assigned to a small-chunk segment; otherwise, it is assigned to a large-chunk segment).
Relationships between the ZapRAID design and the ZNS SSD model.
While our motivating experiments and evaluation are based on the ZN540 model [14], our ZapRAID design does not heavily depend on a specific ZNS SSD model. For example, some of our key design features, such as the group-based data layout and crash recovery management, apply to all ZNS SSD models.
@@
Our ZapRAID design is also applicable to different zone capacities of ZNS SSDs. For smaller zone capacities, the data and footer regions become smaller, yet the group-based data layout remains applicable. For example, when the zone capacity is 96 MiB [62] (or equivalently, 24,576 blocks), for 4-KiB chunks, the header, data, and footer regions occupy 1 block, 24,455 blocks, and 120 blocks, respectively (based on the calculation of §3.1). If the stripe group size is 256, there are still many stripe groups (i.e., 96 groups) for intra-zone parallelism.
Note that the ZN540 model has been widely studied by a large body of prior work in the literature (e.g., [20, 45, 49, 69, 32]), and hence is representative for ZNS SSD research.
@@ 4 Implementation
Dispatch handler.
The dispatch handler is responsible for processing write requests and updating the header and footer regions. For write requests, the dispatch handler assigns the written blocks to segments and submits the I/O events for the written blocks to the device I/O handler. If a block fills a stripe in a segment, the dispatch handler also sends a parity generation event to the encoding handler, which computes parity chunks and submits the I/O events for the parity chunks to the device I/O handler. Recall that the writes of the first stripe of a stripe group need to wait for the previous stripe group to be completely persisted (§3.5); in this case, the dispatch handler will suspend the request and retry later until the segment state is updated by the segment state tracking handler (see below).
@@
Device I/O handler.
The device I/O handler is responsible for submitting block I/O requests to the drives. It receives I/O events from other handlers and submits the corresponding I/O requests to the drives. It is also responsible for checking the completion of each block I/O request from the drives and notifying the completion handler (see below).
Completion handler.
The completion handler tracks the completion state of each I/O request. It receives the notification of completion of a block I/O event from the device I/O handler. If all the I/O events for a request are completed, the completion handler notifies the upper-layer application. The application can register a callback function for each request. For write requests, the completion handler first sends an L2P table update event to the indexing handler, which triggers the callback function after updating the L2P table. For read requests, the completion handler directly triggers the callback function.
@@
Indexing handler.
The indexing handler is responsible for the L2P table maintenance. It receives L2P table query events from the dispatch handler and returns the PBAs of requested blocks for read requests. It also receives L2P table update events from the completion handler and the cleaning handler. If an L2P table update event is from the completion handler, it acknowledges the write requests to the application on behalf of the completion handler after the L2P table is updated. In addition, the indexing handler is responsible for the offloading of L2P table entries if the L2P table size reaches the memory size limit (§3.1).
Encoding handler.
The encoding handler is responsible for parity generation. When it receives a notification from the dispatch handler that a stripe has $k$ data chunks, it generates parity chunks. It then submits I/O events for the parity chunks to the device I/O handler. We handle parity chunk generation outside of the normal I/O requests to mitigate interference on normal I/O requests.
Segment state tracking handler.
The segment state tracking handler examines whether the system needs to write the header region or the footer region of a segment, or whether it has persisted all stripes in a stripe group (§3.5). It periodically (every 1 $\mu$s in our prototype) examines the segment states. When the segment state needs to be updated, it also sends the corresponding I/O events to the device I/O handler.
Cleaning handler.
The cleaning handler is responsible for reclaiming the space of stale blocks. It periodically (every 1 $\mu$s in our prototype) examines the available space and triggers garbage collection if the available space drops below a fixed threshold. It selects one sealed segment that has the most stale blocks and rewrites its valid blocks by submitting block I/O events to the device I/O handler. Currently, our prototype tracks the validity of all blocks (including both user-written blocks and mapping blocks) for each segment via an in-memory bitmap (whose content can also be derived from both the L2P table for user-written blocks and the LBA-to-PBA table for mapping blocks) and collects the LBAs of valid blocks from the block metadata. When all rewrites are completed, the dispatch handler sends the LBAs of all rewritten blocks, along with their old and new PBAs, to the indexing handler for updating the L2P table.
In this work, we do not further consider sophisticated garbage collection algorithms, which need to address the key design issues of (i) when to trigger garbage collection, (ii) which segment to clean for garbage collection, and (iii) where to rewrite the valid data. We argue that such design issues are orthogonal to our current focus on leveraging Zone Append and Zone Write for a high-performance RAID design. We pose the design of new garbage collection algorithms as a future work.
Thread assignment.
We assign the seven handlers to SPDK threads, which exchange events via the function spdk_thread_send_msg and event polling. Currently, we co-locate the dispatch handler, completion handler, segment state tracking handler, and cleaning handler in the main SPDK thread. We also assign the device I/O handler, indexing handler, and encoding handler separately to three different SPDK threads for parallelism, such that our prototype can saturate the write bandwidth of our testbed.
@@ 5 Evaluation
We evaluate ZapRAID through microbenchmarks, trace-driven experiments, and real-application experiments.
@@ 5.1 Methodology
We use a server that runs Ubuntu 22.04 LTS with Linux kernel 5.15. It has a 16-core Intel Xeon Silver 4215 2.5 GHz CPU and 96 GiB DRAM. It is attached with four 4-TiB Western Digital Ultrastar DC ZN540 ZNS SSDs [14]. Each SSD has 3,690 zones, with a zone capacity of 1,077 MiB each. We format each SSD with a logical block size of 4 KiB and a block metadata size (in the out-of-band area) of 64 bytes. By default, we focus on (3+1)-RAID-5, yet we also consider other RAID schemes and a larger-scale emulated ZNS SSD array based on FEMU [53]5.2).
@@
In our RAIZN-SPDK implementation, we focus on achieving high performance and simplify our implementation by removing the metadata zones reserved for write atomicity and garbage collection. RAIZN-SPDK also reserves two metadata zones for high performance. Specifically, it writes partial parity updates to one of the two metadata zones. When the metadata zone becomes full, RAIZN-SPDK resets the zone and writes the partial parity updates to another metadata zone in parallel. For high performance, RAIZN-SPDK uses Zone Append to write partial parity updates to a metadata zone as described in [45]. When there are multiple open segments, RAIZN-SPDK uses Zone Write to write small chunks to small-chunk segments and large chunks to large-chunk segments as in ZoneWrite-Only, while using Zone Append to write partial parity updates of all segments to a metadata zone.
We also attempt to fairly compare ZapRAID with the conventional RAID array that runs on traditional block-interface SSDs. However, the firmware on our ZN540 SSDs is locked and cannot be directly modified, so we cannot make comparisons on the same hardware as in [45]. Thus, we leverage FEMU-emulated SSDs [53] for comparisons and build the conventional RAID array with mdadm [16] (see Exp#6 for details).

### v3

@@ 3.2 Group-Based Data Layout
Trade-off analysis. The choice of $G$ determines the trade-off between the degree of intra-zone parallelism via Zone Append and the stripe management overhead. A larger $G$ allows more stripes to be issued via Zone Append in parallel, but it also increases the stripe management overhead. In the extremes, when $G=1$, all stripes are issued via Zone Write; when $G=S$, there is only one stripe group in each segment (i.e., the group-based data layout is disabled for Zone Append).
@@
In contrast, the case of $G=S$ (i.e., without the group-based data layout) translates to 19 bits per stripe ID and hence 8.95 GiB of memory for the compact stripe table. Also, its high query overhead significantly increases the degraded read latency (§5.4).
Our prototype rounds up the length of a stripe ID to the nearest byte and represents a stripe ID in $\lceil\lceil\log_{2}G\rceil/8\rceil$ bytes. Representing stripe IDs in units of bytes, instead of bits, simplifies data alignment without compromising our conclusion. Thus, for $G=S$, our prototype consumes 3 bytes per stripe ID, leading to 11.3 GiB of memory for the compact stripe table.
@@ 3.3 Hybrid Data Management
Segment management. Figure 5 shows how ZapRAID manages multiple open small-chunk and large-chunk segments. Let $N_{s}$ and $N_{l}$ be the numbers of open small-chunk segments and open large-chunk segments, respectively (i.e., the total number of open zones in each ZNS SSD is $N_{s}+N_{l}$). Also, let $C_{s}$ and $C_{l}$ be the sizes of a small chunk and a large chunk, respectively. All of $N_{s}$, $N_{l}$, $C_{s}$, and $C_{l}$ are configurable parameters. For example, we can set $C_{s}=8$ KiB for the small chunk size and $C_{l}=16$ KiB for the large chunk size. We can also set $N_{s}>1$ and $N_{l}>1$ to accommodate the real-world workloads that are mixed with small and large writes (e.g., in cloud block storage [41]).
@@
Choosing segments from in-flight stripes. ZapRAID maintains in-flight stripes separately for small and large chunks. It assigns new written blocks to an in-flight stripe for either small or large chunks based on the write request size (see above). When an in-flight stripe is formed, ZapRAID forwards the in-flight stripe to an open segment. Specifically, if the in-flight stripe holds large chunks, ZapRAID selects an open large-chunk segment (out of the $N_{l}$) in a round-robin manner and writes the in-flight stripe to the selected segment via Zone Write. On the other hand, if the in-flight stripe holds small chunks, ZapRAID further considers two cases of $N_{s}$. If $N_{s}=1$, it writes the in-flight stripe to the only open small-chunk segment via Zone Append. If $N_{s}>1$, it selects an idle small-chunk segment (out of the $N_{s}-1$) in a round-robin manner and writes the in-flight stripe to the selected segment via Zone Write; however, if there is no idle small-chunk segment, ZapRAID selects the only small-chunk segment reserved for Zone Append and writes the in-flight stripe to the segment via Zone Append.
@@ 3.4 Crash Consistency
Recovery of the segment table. ZapRAID first identifies the open and sealed segments in the system, so as to recover the segment table. Specifically, for each ZNS SSD, ZapRAID checks the states of all zones. If a zone is open or full, ZapRAID retrieves its first block, which belongs to the header region of a segment (§3.1). ZapRAID then reconstructs the segment table with the information of all possible segments based on the first blocks of all open and full zones.
@@
Recovery of stripes. ZapRAID needs to recover the consistent states of all stripes by ensuring that each stripe has persisted $k$ data chunks and $m$ parity chunks. Specifically, ZapRAID examines whether each in-flight stripe has all its $k+m$ chunks persisted. Under the group-based data layout (note that $G=1$ means the use of Zone Write), ZapRAID ensures that only the stripes in the latest stripe group of an open segment can be the in-flight stripes. It retrieves all the chunks in the latest stripe group and examines the stripe IDs in their block metadata. ZapRAID discards all partially persisted stripes with fewer than $k+m$ persisted chunks. Recall that ZapRAID acknowledges the writes of an in-flight stripe only after persisting all $k+m$ chunks of the stripe (§3.1), so discarding the partially persisted stripes does not incur data loss. Since ZNS SSDs cannot perform in-place writes to replace the partially persisted chunks, ZapRAID rewrites all fully persisted stripes in the segment to a new segment and reclaims the old segment.
Recovery of the L2P table and the compact stripe table. After ensuring the consistency of all stripes, ZapRAID recovers the L2P table and the compact stripe table (the latter is for segments that use Zone Append). Recall that for each sealed segment, ZapRAID stores the block metadata (i.e., the LBA, write timestamp, and stripe ID) of all blocks in the segment in the footer region, while for each open segment, ZapRAID also stores the block metadata in the out-of-band area associated with each block. To recover the L2P table and the compact stripe table, ZapRAID first retrieves the footer regions of all sealed segments and the data regions of all open segments. It examines all the block metadata and reconstructs the LBA-to-PBA mappings in the L2P table and the stripe IDs in the compact stripe table. For the L2P table, if multiple blocks have the same LBAs, ZapRAID keeps the one with the latest write timestamp.
@@ 3.5 Complete Workflows
[…]
@@
Reads. To read a block (identified by an LBA), ZapRAID queries the L2P table for its PBA (i.e., the segment ID, the drive ID, and the offset in the respective zone). Given the segment ID and drive ID, ZapRAID locates the zone from the segment table and retrieves the block from the specified offset of the zone.
Degraded reads. Suppose that ZapRAID performs a degraded read to a lost block. Based on the PBA of the requested block, ZapRAID checks whether the segment applies Zone Append or Zone Write. For Zone Append, ZapRAID queries the compact stripe table to find out the stripe ID of the requested block. Since all available chunks of the same stripe must reside in the same stripe group under the group-based data layout, ZapRAID searches for the offsets of the $k$ available chunks in the stripe group from the compact stripe table; note that the search is efficient due to the limited group size (§5.4). It then reads the available chunks from the other zones, decodes the lost chunk, and retrieves the requested block. For Zone Write, ZapRAID reads the available chunks from other zones at the same offset to recover the chunk and hence the requested block.
Full-drive recovery. When a drive fails, ZapRAID recovers the lost data into a new drive. It first identifies the segments that contain the lost zones in the failed drive by examining the segment-to-zones mappings in the header regions of all stored segments. For each lost zone, ZapRAID retrieves all available zones in the same segment from other available drives into memory, and reconstructs the stripe groups that cover the lost zone. It examines the block metadata of the available chunks and identifies the chunks from the same stripe. It then decodes the lost chunks for each stripe independently. After repairing all the stripes in a segment, ZapRAID writes the recovered zone to the new drive.
@@ 3.6 Discussion
Choice of the write request size threshold in hybrid data management. Recall that ZapRAID compares the size of a write request against a threshold to determine if the write request is assigned to a small-chunk or large-chunk open segment (§3.3). Currently, the threshold is set to $C_{l}$. The choice of the threshold poses different performance trade-offs. A larger threshold implies that more writes will go to small-chunk segments. Full small-chunk stripes are formed faster, and hence small writes are also acknowledged faster (i.e., lower response time). However, the write throughput is lower since writing small-chunk segments is slower than writing large-chunk segments (based on the results in Figure 2 for different write request sizes). For simplicity, we assume that the threshold is equal to $C_{l}$ (i.e., if a write request size is smaller than $C_{l}$, it is assigned to a small-chunk segment; otherwise, it is assigned to a large-chunk segment).
Relationships between the ZapRAID design and the ZNS SSD model. While our motivating experiments and evaluation are based on the ZN540 model [70], our ZapRAID design does not heavily depend on a specific ZNS SSD model. For example, some of our key design features, such as the group-based data layout and crash recovery management, apply to all ZNS SSD models.
@@
Our ZapRAID design is also applicable to different zone capacities of ZNS SSDs. For smaller zone capacities, the data and footer regions become smaller, yet the group-based data layout remains applicable. For example, when the zone capacity is 96 MiB [49] (or equivalently, 24,576 blocks), for 4 KiB chunks, the header, data, and footer regions occupy 1 block, 24,455 blocks, and 120 blocks, respectively (based on the calculation of §3.1). If the stripe group size is 256, there are still many stripe groups (i.e., 96 groups) for intra-zone parallelism.
Note that the ZN540 model has been widely studied by a large body of prior work in the literature (e.g., [4, 31, 35, 60, 16]) and is a representative platform for current ZNS SSD research.
@@ 4 Implementation
Dispatch handler. The dispatch handler is responsible for processing write requests and updating the header and footer regions. For write requests, the dispatch handler assigns the written blocks to segments and submits the I/O events for the written blocks to the device I/O handler. If a block fills a stripe in a segment, the dispatch handler also sends a parity generation event to the encoding handler, which computes parity chunks and submits the I/O events for the parity chunks to the device I/O handler. Recall that the writes of the first stripe of a stripe group need to wait for the previous stripe group to be completely persisted (§3.5); in this case, the dispatch handler will suspend the request and retry later until the segment state is updated by the segment state tracking handler (see below).
@@
Device I/O handler. The device I/O handler is responsible for submitting block I/O requests to the drives. It receives I/O events from other handlers and submits the corresponding I/O requests to the drives. It is also responsible for checking the completion of each block I/O request from the drives and notifying the completion handler (see below).
Completion handler. The completion handler tracks the completion state of each I/O request. It receives the notification of completion of a block I/O event from the device I/O handler. If all the I/O events for a request are completed, the completion handler notifies the upper-layer application. The application can register a callback function for each request. For write requests, the completion handler first sends an L2P table update event to the indexing handler, which triggers the callback function after updating the L2P table. For read requests, the completion handler directly triggers the callback function.
@@
Indexing handler. The indexing handler is responsible for the L2P table maintenance. It receives L2P table query events from the dispatch handler and returns the PBAs of requested blocks for read requests. It also receives L2P table update events from the completion handler and the cleaning handler. If an L2P table update event is from the completion handler, it acknowledges the write requests to the application on behalf of the completion handler after the L2P table is updated. In addition, the indexing handler is responsible for the offloading of L2P table entries if the L2P table size reaches the memory size limit (§3.1).
Encoding handler. The encoding handler is responsible for parity generation. When it receives a notification from the dispatch handler that a stripe has $k$ data chunks, it generates parity chunks. It then submits I/O events for the parity chunks to the device I/O handler. We handle parity chunk generation outside of the normal I/O requests to mitigate interference on normal I/O requests.
[…]
In this work, we do not further consider sophisticated garbage collection algorithms, which need to address the key design issues of (i) when to trigger garbage collection, (ii) which segment to clean for garbage collection, and (iii) where to rewrite the valid data. We argue that such design issues are orthogonal to our current focus on leveraging Zone Append and Zone Write for a high-performance RAID design. We pose the design of new garbage collection algorithms as future work.
Thread assignment. We assign the seven handlers to SPDK threads, which exchange events via spdk_thread_send_msg and event polling. Currently, we co-locate the dispatch handler, completion handler, segment state tracking handler, and cleaning handler in the main SPDK thread. We also assign the device I/O handler, indexing handler, and encoding handler separately to three different SPDK threads for parallelism, such that our prototype can saturate the write bandwidth of our testbed.
@@ 5 Evaluation
We evaluate ZapRAID through real-application experiments, trace-driven experiments, and microbenchmarks.
@@ 5.1 Methodology
We use a server that runs Ubuntu 22.04 LTS with Linux kernel 5.15. It has a 16-core Intel Xeon Silver 4215 2.5 GHz CPU and 96 GiB DRAM. It is attached with four 4 TiB Western Digital Ultrastar DC ZN540 ZNS SSDs [70]. Each SSD has 3,690 zones, with a zone capacity of 1,077 MiB each. We format each SSD with a logical block size of 4 KiB and a block metadata size (in the out-of-band area) of 64 bytes. By default, we focus on (3+1)-RAID-5, yet we also consider other RAID schemes and a larger-scale emulated ZNS SSD array based on FEMU [39]5.4).
@@
In our RAIZN-SPDK implementation, we focus on achieving high performance. We simplify our implementation by removing the metadata zones reserved for write atomicity and garbage collection. RAIZN-SPDK also reserves two metadata zones for high performance. Specifically, it writes partial parity updates to one of the two metadata zones. When the metadata zone becomes full, RAIZN-SPDK resets the zone and writes the partial parity updates to another metadata zone in parallel. For high performance, RAIZN-SPDK uses Zone Append to write partial parity updates to a metadata zone as described in [31]. When there are multiple open segments, RAIZN-SPDK uses Zone Write to write small chunks to small-chunk segments and large chunks to large-chunk segments, as in ZoneWrite-Only, while using Zone Append to write partial parity updates of all segments to a metadata zone.
We also attempt to fairly compare ZapRAID with the conventional RAID array that runs on traditional block-interface SSDs. However, the firmware on our ZN540 SSDs is locked and cannot be directly modified, so we cannot make comparisons on the same hardware as in [31]. Thus, we leverage FEMU-emulated SSDs [39] for comparisons and build the conventional RAID array with mdadm [47] (see Exp#10 for details).